Show the effect on the demand, the supply, the equillibrium price, and the equilibrium quantity of each of the following events on the market for beef.


a. Case 1: The salaries of workers increases in the beef market

b. Case 2: Outbreak of mouth and food disease in the region.

c. Case 3: New information association heart diseases to red meat (beef)

d. Case 4: The increase in the price of chickens

e. Case 5: Government introduce subsidy on beef.



1.1 Suppose a firm sells 20000 units when the price is $16 but sells 30000 units when the price falls to $14.


a. Calculate the percentage change in the quality sold over this price range using the midpoint formula.

b. Calculate the percentage change in the price using the midpoint formula.

c. Find the price elasticity of demand over this range of prices.

d. State whether demand is elastic or inelastic over this range..


1
Expert's answer
2020-07-28T10:49:19-0400

a. Case 1: The salaries of workers increases in the beef market



The proposal is changing. In the chart, the initial offer is shown with a thick green line and the resulting offer is shown with a thin green line. Price increases, less meat is sold.


b. Case 2: Outbreak of mouth and food disease in the region.



Demand is changing. In the chart, the initial demand is shown with a blue line and the resulting demand is shown with a red line. Prices are falling, less meat is sold.

c. Case 3: New information association heart diseases to red meat (beef)



Demand is changing. In the chart, initial demand is shown with a thin blue line, and the resulting demand is shown with a thick blue line. Prices are falling, less meat is sold.

d. Case 4: The increase in the price of chickens



Demand is changing. These goods are interchangeable. In the chart, the initial demand is shown as a red line, and the resulting demand is shown as a blue line. Price increases, more beef is sold.

e. Case 5: Government introduce subsidy on beef.



The offer is changing. In the chart, the initial supply is shown with a green line and the resulting supply is shown with an orange line. The price goes down, more beef is sold.


1.1 Suppose a firm sells 20000 units when the price is $16 but sells 30000 units when the price falls to $14.


a. Calculate the percentage change in the quality sold over this price range using the midpoint formula.


"\varDelta Q= \frac {Q_2-Q_1}{(Q_2+Q_1)/2} \times 100=40"

b. Calculate the percentage change in the price using the midpoint formula.


"\varDelta p= \frac {p_2-p_1}{(p_2+p_1)/2} \times 100=13"


c. Find the price elasticity of demand over this range of prices.


"E=\frac {\varDelta Q}{ \varDelta p}=3.08"

d. State whether demand is elastic or inelastic over this range..

The demand is elastic, since the obtained coefficient is greater than 1.